OpenStudy (anonymous):

what is the solution(s) for 3^x=5x

OpenStudy (anonymous):

is it \(3^x=5x\) really?

OpenStudy (anonymous):

i would perhaps graph of use wolfram you are not going to solve this one with algebra

OpenStudy (anonymous):

yeah, its on my worksheet. and im in algebra 1..

OpenStudy (anonymous):

it really says solve three to the power of x equals five times x ???

OpenStudy (anonymous):

yep, and theres nothing to plug in for x.

OpenStudy (anonymous):

not maybe \(x^3=5x\)?

OpenStudy (anonymous):

can you attach it in some way, because there is no algebra to do this one

OpenStudy (anonymous):

promise. I'm looking at it right now.

OpenStudy (anonymous):

I can't. I guess I;ll just ask in class tomorrow...haha.

OpenStudy (anonymous):

i believe you i am not sure what you are supposed to do exactly to solve this

OpenStudy (anonymous):

x^2 * x = 5x x^2 = 5 so, what are the 3 possible values of x....

OpenStudy (anonymous):

@jon.stromer.galley it is not \(x^3\) it is \(3^x\) that is way different

OpenStudy (anonymous):

you can check an answer here by clicking on "approximate form" but how to get it i would have no idea http://www.wolframalpha.com/input/?i=3^x%3D5x

OpenStudy (anonymous):

oops, sorry I mis-read it. Yes 3^x is a whole different ball of wax

OpenStudy (anonymous):

and if there is an algebra one way to solve this, i will eat my bicycle i wonder what the teacher will say...

OpenStudy (anonymous):

hahaha. thank y'all for trying to help. If I remember, I will log back on and tell you what she says.

OpenStudy (anonymous):

If that x was away from the 5, it would be much easier.